# ChartAnchor: Chart Grounding with Structural-Semantic Fidelity and Data Recovery

### v1

@@ 4. Evaluation Metrics
To assess prediction quality, we adopt a matching scheme inspired by the Structuring Chart-oriented Representation Metric (SCRM) (xia2023structchart). For each predicted tuple $p$ and ground-truth tuple $q$, we compare corresponding fields $p_{i}$ and $q_{i}$, where $i=0,1,\dots,n{-}1$ and $n$ is the tuple length. String fields are evaluated using edit distance $J(p_{i},q_{i})$, and numerical fields using relative error $e(p_{i},q_{i})$. A tuple is considered correct only if all fields satisfy their respective tolerance, with three tolerance levels: strict ($J\leq 0$ or $e\leq 0$), slight ($J\leq 3$ or $e\leq 0.05$), and high ($J\leq 5$ or $e\leq 0.10$).
We report precision, recall, F1, and Intersection over Union (IoU) under a given tolerance level as:
$P_{tol}=\frac{|M|}{|\text{Pred}|},\quad R_{tol}=\frac{|M|}{|\text{GT}|},\quad$
$\text{F1}_{tol}=\frac{2\cdot P_{tol}\cdot R_{tol}}{P_{tol}+R_{tol}},\quad\text{IoU}_{tol}=\frac{|M|}{|\text{Pred}|+|\text{GT}|-|M|}$
Here, $|M|$ is the number of matched tuples under tolerance, and $|\text{Pred}|$, $|\text{GT}|$ are the counts of predicted and ground truth tuples, respectively. This metric provides a principled and fine-grained assessment of semantic alignment and numerical fidelity, offering a robust measure of structured data reconstruction across both tasks.
@@
Perceptual Similarity. To bridge the gap between syntactic correctness and human-like perception, We employ CLIPScore to evaluate image semantic consistency through embedding alignment. This complements structured evaluations by quantifying conceptual fidelity grounded in visual cognition.
5. Evaluation
@@
In this section, we present a comprehensive evaluation of various MLLMs using our ChartAnchor benchmark. Our assessment includes both state-of-the-art proprietary (closed-source) models and representative open-source MLLMs, offering a holistic comparison of their performance.
@@
GPT4o 97.12 10.74 8.44 8.65 37.71 23.30 24.69 60.42 35.44 37.72
Claude-3-7-Sonnet 94.85 12.75 11.02 10.86 44.64 29.43 30.52 58.12 37.11 38.62
@@
InternVL3-2B 55.16 5.84 3.86 4.08 22.18 10.88 11.98 51.04 18.29 21.19
Qwen2.5-VL-3B-Instruct 92.18 9.11 6.13 6.51 36.70 17.28 19.12 61.01 27.43 30.73
Gemma-3-4B-it 96.23 7.35 5.38 5.52 32.05 18.33 19.72 48.89 26.49 28.60
DeepSeek-VL-7B 98.34 1.72 0.36 0.49 5.47 0.90 1.33 26.95 4.79 7.02
LLaVA-v1.6-Mistral-7B 76.11 4.57 1.66 2.02 22.22 6.33 7.72 44.38 12.15 14.89
Qwen2.5-VL-7B-Instruct 95.05 12.17 8.18 8.67 48.45 23.75 26.35 69.58 30.96 34.90
MiniCPM-V-2.6-8B 79.66 5.07 3.03 3.30 26.80 11.07 12.51 53.36 18.98 22.09
InternVL3-9B 88.43 11.42 8.79 8.95 44.54 25.36 26.98 63.49 34.47 36.96
GLM-4V-9B 91.93 7.14 4.86 5.12 27.42 13.17 14.50 59.50 24.64 27.70
InternVL3-14B 66.18 8.19 7.47 7.28 29.27 22.30 22.81 40.90 29.67 30.49
CogVLM2-Llama3-Chat-19B 78.41 2.43 0.87 1.06 10.28 3.92 4.56 25.84 8.37 10.36
Qwen2.5-VL-32B-Instruct 97.83 13.05 9.87 10.10 48.65 28.94 30.99 68.14 37.97 41.04
@@
Table 5. Comparison of model performance on the controlled chart-to-table task. Pass Rate indicates the proportion of examples for which the model produced a parsable table. P, R, and F1 denote precision, recall, and F1-score, respectively. Colored markers represent different tolerance levels:   for strict,   for slight, and   for high.
5.1. Benchmarking Models
We evaluate 14 MLLMs. For proprietary models, we include two state-of-the-art enterprise-level models: GPT-4o (openai2024) and Claude-3-7-Sonnet (claude-sonnet), which represent the forefront of commercially deployed multimodal AI systems. On the open-source side, we consider a diverse set of architectures, primarily vision-language models (VLMs), across various parameter scales. In the compact category (under 7B parameters), we evaluate InternVL3-2B(zhu2025internvl3), Qwen2.5-VL-3B-Instruct(yang2024qwen2) and Gemma-3-4B-it(team2025gemma) for their efficiency and lightweight deployment potential. The mid-range segment (7–14B parameters) includes GLM-4V-9B(glm2024chatglm), InternVL3-9B(zhu2025internvl3), DeepSeek-VL-7B-chat(lu2024deepseek), LLaVA-v1.6-Mistral-7B-HF(li2024llava), MiniCPM-V-2.6-8B(yao2024minicpm), and Qwen2.5-VL-7B-Instruct(yang2024qwen2), which balance computational cost and multimodal reasoning capabilities through advanced attention mechanisms. At the large scale (19–32B parameters), we assess CogVLM2-Llama3-Chat-19B(hong2024cogvlm2), InternVL3-14B(zhu2025internvl3), and Qwen2.5-VL-32B-Instruct(yang2024qwen2), all of which leverage powerful visual encoders and extended context handling to tackle complex vision-language understanding tasks.
5.2. Experimental Setting and Resource
For all models, we set the temperature to $\tau=0.1$ and use top-$p$ sampling with $p=0.95$ for decoding. The maximum generation length is capped at 4096 tokens. For open-weight models, we adopt bfloat16 precision during inference. All experiments are conducted on A100 40G GPUs. The prompts used for each task are detailed in Appendix E.
5.3. Key Insights from Results
Tab. 4 reports the performance of the evaluated models on the Chart-to-Code task, whereas Tab. 5 presents their results on the controlled Chart-to-Table reconstruction task. Based on these results, we derive several key insights, with further discussion provided in the Appendix E.
@@ 5.3. Key Insights from Results
Figure 5. Comparison of Data Attribute Prediction Across Chart Types for Different Models
1) GPT-4o demonstrates the best overall performance on the Chart-to-Code task, while Claude-3-7-Sonnet leads in the Controlled Chart-to-Table task. GPT-4o’s strength lies in its superior visual reasoning and structural consistency, yielding high pass rates and fidelity scores in code generation; meanwhile, Claude-3-7-Sonnet shows stronger table reconstruction capabilities, likely due to its alignment with structured data tasks—both clearly outperform open-source models.
2) Proprietary models lead in overall chart grounding, but open-source models are closing the gap in specific sub-tasks.GPT-4o achieves the highest overall score in chart-to-code generation (63.02) and strong F1 scores in chart-to-table reconstruction (F1-High: 37.72), outperforming all open models. However, Qwen2.5-VL-32B-Instruct surpasses GPT-4o in F1-High (41.04).
3) Model scale correlates with performance within model families, but architecture and training dominate cross-family comparisons.Overall scores increase with scale in Qwen2.5-VL and InternVL3 series(e.g., Qwen-VL: 3B–28.63, 7B–43.68, 32B–48.76; InternVL3: 2B–17.62, 9B–45.29, 14B–48.76). In contrast, CogVLM2-19B scores only 4.35 overall, suggesting architectural design and training alignment are more decisive than size alone.
4) Smaller models are competitive when aligned with task-specific objectives. Gemma-3-4B-it performs comparably to or better than some 7B+ models in layout accuracy and slight tolerance F1, showing that targeted training can offset limited scale, especially in lightweight deployment settings.
@@
5) Chart-to-code generation reveals consistent weaknesses in color replication and data fidelity. Even top models (e.g., GPT-4o, InternVL3-14B) achieve lower scores in Color and Data Fidelity compared to structural elements like Title or Layout, indicating that fine-grained visual detail understanding remains a major bottleneck in generative chart reasoning.
6) Tolerance analysis highlights limitations in precise numerical reasoning. Across models, strict-F1 scores are substantially lower than those under slight or high tolerance, confirming that while macro-structure comprehension is relatively robust, exact data recovery—especially for decimals and axis-scaled values—remains unreliable.
7) Best Performance per Size. Among open-source models, Qwen2.5-VL-32B stands out as the top large model, coming close to GPT-4 in several areas. But for lower-budget scenarios, the 7B–14B tier models like InternVL3-9B and Qwen2.5-VL-7B offer strong performance at a fraction of the size. They both surpass 43 overall points, which is impressive relative to many other models under 10B. Meanwhile, certain 7B models not specifically tuned for this domain (like LLaVA 1.6 or DeepSeek) lag far behind. The comparison suggests that investing in domain-specific fine-tuning yields more benefit than simply adding parameters. A well-trained medium model can beat an untuned large model by a wide margin. This parameter efficiency perspective is crucial for real-world use, where deploying a smaller model that achieves, say, 80% of GPT-4’s performance could be far more practical.
8) The CLIP-based semantic alignment score tends to correlate positively with visual structure metrics like chart type and layout. For example, models with high layout alignment (GPT-4 at 80.7%, Claude-3-7-Sonnet at 76.5%) also show high CLIP scores (GPT-4 86.9, Claude-3-7-Sonnet 94.1) in Chart-to-Code. This suggests CLIP is a reliable proxy for detecting major structural errors—when models misidentify layout or chart type, CLIP similarity typically drops. While not sufficient for fine-grained correctness, CLIP complements structural metrics by capturing overall visual alignment.
5.4. Type Analysis
Fig.  5 presents the accuracy of the "data" attribute prediction across 30 chart types on the Chart-to-Code task, categorized into seven semantic groups in the context of the chart2code task. Among basic 2D Cartesian plots (e.g., bar, line, scatter), GPT-4o consistently outperforms other models, achieving over 0.8 accuracy in pie and areachart, while both Gemma3-4b and InternVL3-14b demonstrate stable but lower performance.
For polar plots such as barpolar and funnelarea, all models exhibit a marked performance drop, reflecting challenges in handling angular and radial encoding. In the 3D and high-dimensional category (scatter3d, surface, cone), accuracy further declines across the board, underscoring the difficulty of abstract spatial structures.
Matrix-style charts like heatmap and his2dcontour show modest improvements for GPT-4o, but still lag due to their dense grid-based semantics. In multidimensional projections (parcoords, scatterternary), prediction accuracy is near zero, suggesting current models lack sufficient capability to model complex mapping relationships.
Hierarchical and network-style charts (treemap, sunburst, sankey) reveal slight gains for GPT-4o, while others fail to generalize. Notably, in specialized financial plots (candlestick, ohlc, waterfall), GPT-4o again achieves high scores, indicating its better domain transfer ability. Overall, while GPT-4o demonstrates broad robustness, notable gaps remain in its ability to handle structurally complex or semantically specialized chart types, highlighting opportunities for future research.
6. Conclusion
@@
In this work, we introduced ChartAnchor, a comprehensive benchmark for rigorous evaluating chart grounding in MLLMs. Unlike prior benchmarks that evaluate isolated tasks or focus on limited chart types, ChartAnchor provides a unified framework for both chart-to-code generation and controlled chart-to-table reconstruction, enabling holistic assessment of visual, structural, and numerical fidelity. It contains 8,068 chart–table–code triplets across 30 diverse chart types and two plotting libraries, capturing the complexity and heterogeneity of real-world visualizations. We further proposed a multi-level evaluation protocol that includes functional validity, visual consistency, data fidelity and perceptual similarity, and data accuracy. Empirical results across 14 leading MLLMs reveal that while state-of-the-art models such as GPT-4o demonstrate strong performance in structural parsing and layout reproduction, they still struggle with fine-grained data recovery and color fidelity. These findings underscore the pressing need for models that integrate symbolic reasoning and precise visual understanding. We hope ChartAnchor catalyzes future research on robust, semantically grounded chart comprehension, particularly in domains like science, finance, and policy where both interpretability and data integrity are paramount.
7. Limitations and Future Work
While the current version of ChartAnchor focuses on static chart understanding, we recognize that real-world visualizations increasingly involve interactive or dynamic elements—such as drill-down plots, animated data transitions, or multi-view dashboards. These charts introduce additional layers of complexity, including temporal structure, user interaction states, and evolving visual contexts.
In future work, ChartAnchor can be extended to incorporate interactive and time-varying visualizations. Specifically, benchmarks could be designed to simulate interaction flows and temporal data changes, enabling the evaluation of model capabilities in handling dynamic semantics and multi-state rendering. Such an extension would further broaden the applicability of ChartAnchor to real-world analytical environments.
@@ Appendix A Detailed Data Collection and Processing Procedures
A.1. Data Sources and Corpus Composition
@@
Table 6. Number of samples per chart type in the newly curated dataset.
@@
A.2. Detailed Code Generation and Augmentation for Existing Datasets

### v3

@@ 4 Evaluation Metrics
To assess prediction quality, we adopt a matching scheme inspired by the Structuring Chart-oriented Representation Metric (SCRM) (Xia et al., 2023). For each predicted tuple $p$ and ground-truth tuple $q$, we compare corresponding fields $p_{i}$ and $q_{i}$, where $i=0,1,\dots,n{-}1$ and $n$ is the tuple length. String fields are evaluated using edit distance $J(p_{i},q_{i})$, and numerical fields using relative error $e(p_{i},q_{i})$. A tuple is considered correct only if all fields satisfy their respective tolerance, with three tolerance levels: strict ($J\leq 0$ or $e\leq 0$), slight ($J\leq 3$ or $e\leq 0.05$), and high ($J\leq 5$ or $e\leq 0.10$).
We evaluate structural reconstruction using Precision ($P=n_{m}/n_{p}$), Recall ($R=n_{m}/n_{gt}$), F1 score ($2PR/(P+R)$), and IoU ($n_{m}/(n_{p}+n_{gt}-n_{m})$). Here, $n_{m}$, $n_{p}$, and $n_{gt}$ denote the counts of matched, predicted, and ground-truth tuples, respectively. These metrics provide a principled assessment of semantic alignment and numerical fidelity.
@@
Perceptual Similarity. To bridge the gap between syntax and human perception, we employ CLIPScore to measure semantic consistency via embedding alignment. This complements structured metrics by quantifying conceptual fidelity grounded in visual cognition.
5 Evaluation
@@
This section benchmarks diverse proprietary and open-source MLLMs using ChartAnchor, offering a holistic performance comparison.
5.1 Evaluated Models and Implementation
We benchmark 24 MLLMs across two categories. For the open-source models, total parameters ranging from 2B to 235B, including: Qwen2.5-VL series, Qwen3-VL series, InternVL3 series, DeepSeek-VL series, GLM series, and MiMo-VL-7B variants, alongside Gemma-3-4B-it, LLaVA-v1.6-Mistral-7B, MiniCPM-V-2.6-8B, and CogVLM2-Llama3-Chat-19B. For proprietary models, we include GPT-4o, GPT-5, Claude-3.7-Sonnet, Claude-4.5-Sonnet, and Gemini-3-Pro. For reproducibility, detailed experimental settings and the full list of prompts are relegated to Appendix E.
@@ 5.1 Evaluated Models and Implementation
GPT-4o (OpenAI, 2024) 97.12 10.74 8.44 8.65 37.71 23.30 24.69 60.42 35.44 37.72
GPT-5 (OpenAI, 2025) 99.94 19.62 17.88 17.64 61.41 44.63 45.89 77.48 53.79 55.36
Claude-3-7-Sonnet (Anthropic, 2025a) 94.85 12.75 11.02 10.86 44.64 29.43 30.52 58.12 37.11 38.62
Claude-4-5-Sonnet (Anthropic, 2025a) 98.66 14.89 14.46 13.65 43.94 40.22 38.29 54.87 49.6 47.11
Gemini-3-Pro (Comanici et al., 2025) 95.19 19.96 18.03 18.15 57.78 45.18 46.48 70.61 52.93 54.72
@@
InternVL3-2B (Zhu et al., 2025) 55.16 5.84 3.86 4.08 22.18 10.88 11.98 51.04 18.29 21.19
Qwen2.5-VL-3B-Instruct (Yang et al., 2024a) 92.18 9.11 6.13 6.51 36.70 17.28 19.12 61.01 27.43 30.73
Gemma-3-4B-it (Team et al., 2025c) 96.23 7.35 5.38 5.52 32.05 18.33 19.72 48.89 26.49 28.60
DeepSeek-VL2-27B-A4B (Wu et al., 2024b) 75.18 3.81 3.03 3.08 17.33 12.15 12.76 24.83 15.65 16.76
DeepSeek-VL-7B (Lu et al., 2024) 98.34 1.72 0.36 0.49 5.47 0.90 1.33 26.95 4.79 7.02
LLaVA-v1.6-Mistral-7B (Li et al., 2024a) 76.11 4.57 1.66 2.02 22.22 6.33 7.72 44.38 12.15 14.89
Qwen2.5-VL-7B-Instruct (Yang et al., 2024a) 95.05 12.17 8.18 8.67 48.45 23.75 26.35 69.58 30.96 34.90
MiMo-VL-7B-SFT (Team et al., 2025b) 97.5 12.38 9.71 9.91 47.48 28.72 30.5 70.72 38.2 41.09
MiMo-VL-7B-RL (Team et al., 2025b) 97.36 12.55 9.86 10.07 47.63 28.92 30.64 68.91 37.96 40.64
MiniCPM-V-2.6-8B (Yao et al., 2024) 79.66 5.07 3.03 3.30 26.80 11.07 12.51 53.36 18.98 22.09
Qwen3-VL-8B-Instruct (Bai et al., 2025) 92.76 13.18 10.02 10.31 51 29.44 31.63 69.2 36.7 39.97
InternVL3-9B (Zhu et al., 2025) 88.43 11.42 8.79 8.95 44.54 25.36 26.98 63.49 34.47 36.96
GLM-4V-9B (GLM et al., 2024b) 91.93 7.14 4.86 5.12 27.42 13.17 14.50 59.50 24.64 27.70
InternVL3-14B (Zhu et al., 2025) 66.18 8.19 7.47 7.28 29.27 22.30 22.81 40.90 29.67 30.49
CogVLM2-Llama3-Chat-19B (Hong et al., 2024) 78.41 2.43 0.87 1.06 10.28 3.92 4.56 25.84 8.37 10.36
Qwen2.5-VL-32B-Instruct (Yang et al., 2024a) 97.83 13.05 9.87 10.10 48.65 28.94 30.99 68.14 37.97 41.04
Qwen3-VL-32B-Instruct (Bai et al., 2025) 97.56 16.04 12.92 13.15 55.91 35.9 37.86 72.7 43.53 46.37
GLM-4.5V-106B (Team et al., 2025a) 99.8 14.13 11.71 11.83 52.7 34.48 36.12 71.98 43.49 45.87
Qwen3-VL-235B-A22B-Instruct (Bai et al., 2025) 99.58 17.59 13.75 14.19 63.76 37.97 40.57 85.25 45.7 49.45
@@
Table 5: Comparison of model performance on the controlled chart-to-table task. Pass Rate indicates the proportion of examples for which the model produced a parsable table. P, R, and F1 denote precision, recall, and F1-score, respectively. Colored markers represent different tolerance levels:   for strict,   for slight, and   for high.
@@
Figure 5: Comparison of Data Attribute Prediction Across Chart Types for Different Models
5.2 Key Insights from Results
Tab. 4 and 5 present the results for the Chart-to-Code and controlled Chart-to-Table reconstruction tasks, respectively. Key insights are discussed below, with further details in Appendix E.
1) GPT-5 establishes a new state-of-the-art across both tasks. GPT-5 achieves the highest overall score in chart-to-code generation (72.93) and strong F1 scores in chart-to-table reconstruction (F1-High: 55.36). Its superior visual reasoning capabilities are evidenced by exceptional pass rates and leading scores in both Visual Structure Consistency and Data Fidelity. While open-source models exhibit recent advancements, a distinct performance gap remains, particularly regarding generation accuracy and structural robustness compared to the proprietary baseline.
2) The "Thinking" Paradox: Reasoning enhances structural alignment but risks amplifying perceptual errors. Models equipped with "thinking" modes (e.g., GPT-5, Gemini 3) excel in layout alignment and chart type recognition by logically deducing structural composition. However, this capabilities does not guarantee improvements in data or color fidelity; in fact, reasoning chains may inadvertently amplify hallucinations by logically rationalizing initial perception errors. Ultimately, "thinking" upgrades logical planning rather than low-level visual acuity.
3) On-Policy RL catalyzes structural reasoning and code synthesis. Relative to MiMo-VL-7B-SFT, MiMo-VL-7B-RL achieves a substantial qualitative leap in Chart-to-Code generation, elevating the pass rate from 35.56 to 41.51 and overall score from 23.30 to 28.10. Driven by the Mixed On-policy RL (MORL) framework, the model transcends the surface-level mimicry characteristic of SFT to internalize visual logical hierarchies. This advancement is evidenced by superior Visual Structure Consistency—where RL effectively aligns COT reasoning with the rigorous global topological constraints of code syntax.
@@ 5.2 Key Insights from Results
4) Chart-to-code generation reveals consistent weaknesses in color replication and data fidelity. Even top models (e.g., GPT-5) achieve lower scores in Color and Data Fidelity compared to structural elements like Title or Layout, indicating that fine-grained visual detail understanding remains a major bottleneck in generative chart reasoning.
5.3 Type Analysis
Figure 5 shows a comparative analysis of data attribute prediction accuracy across chart types in the Chart-to-Code task. Overall, GPT-5 establishes a performance ceiling, particularly in 2D Cartesian and Finance/Series categories, while Qwen3-VL remains a competitive baseline, often matching GPT-5 in fundamental types like pie and histogram.
Crucially, we observe a substantial Geometric Reasoning Gap. Performance for all models degrades sharply in Polar and 3D categories. While models handle rectilinear layouts effectively, performance collapses on their radial or multi-dimensional counterparts. This indicates that current VLMs rely heavily on visual pattern recognition rather than intrinsic geometric reasoning. They struggle to map pixel-space information to non-orthogonal coordinate systems or to resolve spatial occlusions in 3D projections, highlighting a lack of spatial inductive bias for complex topologies.
Furthermore, all models, including GPT-5, exhibit a Struggle with Statistical Abstraction. They underperform on Matrix and abstract statistical charts (e.g., box, violin, heatmap). Unlike scatter plots that map raw data points directly, these charts require interpreting aggregated statistical distributions or dense grid encodings. This implies that chart-to-code capabilities are currently constrained by the models’ limited ability to deconstruct high information density and interpret abstract visual summaries.
5.4 Alignment with Human Evaluation
Metric Acc (%) Kendall’s $\tau$
Visual Structure Consistency 88.0 0.76
Perceptual Quality 84.0 0.68
Semantic Data Fidelity 85.0 0.71
Table 6: Human alignment analysis.
To verify whether our evaluation metrics reflect human preferences, we conduct a pairwise comparison study. We sample 300 chart pairs generated by different models from the same ground-truth input and ask domain experts to indicate their preference along Visual Structure Consistency, Semantic Data Fidelity, and Perceptual Quality. As shown in Tab. 6, both Consistency Accuracy and Kendall’s $\tau$ demonstrate strong agreement between metric-induced rankings and human judgments. Experimental setup and annotation details are provided in Appendix E.5.
6 Conclusion
@@
In this work, we introduce ChartAnchor, a comprehensive benchmark for evaluating chart grounding. Unlike prior tasks, ChartAnchor unifies chart-to-code generation and controlled table reconstruction to rigorously assess visual, structural, and numerical fidelity. The dataset encompasses 8,068 chart–table–code triplets spanning 30 diverse types and multiple plotting libraries. We further proposed a multi-level evaluation protocol covering functional validity, visual consistency, data accuracy and perceptual similarity. Experiments across 24 MLLMs reveal that even the performance leader, GPT-5, continues to struggle with fine-grained data recovery and visual structure consistency. These findings underscore the need for integrating symbolic reasoning with visual precision, and we hope ChartAnchor catalyzes research on robust chart comprehension in critical domains.
7 Limitations and Future Work
While ChartAnchor currently targets static charts, real-world visualizations increasingly leverage interactive and dynamic components. Future work aims to extend the benchmark to encompass a broader spectrum of dynamics, such as drill-down plots, animated data transitions, and multi-view dashboards. This expansion enables the evaluation of model proficiency in handling dynamic semantics and multi-state rendering, significantly broadening the benchmark’s relevance to realistic analytical environments.
@@ Appendix A Detailed Data Collection and Processing Procedures
A.1 Data Sources and Corpus Composition
@@
Table 7: Number of samples per chart type in the newly curated dataset.
@@
A.2 Detailed Code Generation and Augmentation for Existing Datasets